Please help!
Solve by substitution method.
7m+n=7
m-8n=58
m=7-7m
m-8(7-7m)=58
m-56+56m=58
?=114

Answer by Alan3354(69443)

Please help!
Solve by substitution method.
7m+n=7
m-8n=58
--------
n=7-7m
m-8(7-7m)=58
m-56+56m=58
57m = 114
m = 2
-----
n = 7-7m
n = -7